# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@
Figure 4: Repeated ancilla-based quantum logic spectroscopy (QLS). a, We consider three ensembles of atoms: reference atoms (R; red), clock atoms (C; purple), and ancillas (A; blue). An entangling gate maps the state of clock atoms to the state of ancilla atoms before mid-circuit readout (MCR) of the ancilla state. b, Logic spectroscopy via ancilla atoms. A Ramsey interrogation is performed in parallel on reference atoms (R) and clock atoms (C). A clock-ancilla anti-CNOT gate maps the state onto ancilla atoms (A) for readout. c, Ancilla-based $\hat{X}$ measurement. A non-destructive measurement of the state $|\psi\rangle_{C}$ is realized by shelving clock qubits ($S$ operation) followed by ancilla MCR. We show that following unshelving ($S^{\dagger}$), the state of clock qubits is projected to a well-defined state in the measurement basis, conditioned on the ancilla state. d, $\rm{N_{cyc}}$ repeated rounds of QLS with minimal (2.9 ms) dead time, enabled by ancilla replacement via array reconfiguration. Each round $j$ consists of a fixed interrogation time $T_{j}$ followed by ancilla-based $\hat{X}$ measurement before a final round where the clock qubit is directly measured. e, We find improved contrast in the final measurement round when post-selecting on ancilla results (see text) in all previous rounds (purple markers) as compared to reference clock atoms which are protected during MCR but do not interact with ancillas (red markers).
@@
Having used entanglement to prepare metrological probe states, we now turn to use it to read out the resulting state. Specifically, we demonstrate a method based on mapping quantum information from a set of sensor qubits to another set of ancillary qubits, which are measured while the sensor qubits are quickly available for sensing once more. Such repeated, non-destructive, ancilla-based measurements, which are central to quantum error detection and correction 39, as well as schemes benefiting from such measurements, have remained largely unexplored with neutral atoms.
@@
To efficiently realize ancilla-based readout in atomic systems, information-carrying data qubits must first be protected from global imaging light and their coherence must be preserved during mid-circuit readout (MCR) of ancilla qubits. Such protection can be realized by using dual-species arrays 52, cavity-assisted readout 53, a separate readout zone 24, or shelving to states dark to imaging light 54, 55, 56. In our system, we take the last approach and have recently realized MCR 44 by shelving optical qubits into superpositions of motional states which are long-lived and dark to imaging light.
@@
By performing high-fidelity entangling gates between information-carrying (clock) qubits and ancilla qubits before MCR (Fig. 4a), we can realize a form of quantum logic spectroscopy (QLS) with neutral atom optical clock qubits, as originally introduced for mixed-species ion traps 20. In QLS, a non-destructive, indirect readout of an atom is realized by mapping its state onto an ancillary atom, followed by a readout of the ancilla. In our case, we use local control to place certain atoms in superposition states sensitive to the clock laser detuning $\Delta$, while ancilla atoms remain in the insensitive state $|1\rangle$. A Ramsey sequence maps $\Delta$ into population imbalance of clock atoms. A subsequent clock-ancilla anti-CNOT gate then maps this signal onto the ancilla state (Fig. 4b).
@@
To realize a non-destructive ancilla-based $\hat{X}$ measurement, we then shelve clock qubits (using motional state shelving 44) and read out the ancilla state. To this end, we ramp up the depth of ancilla traps by a factor of 5 (leaving traps with clock atoms unchanged) such that they are non-resonant with the global motional shelving drive. Following shelving of clock qubits, we perform ancilla MCR which has two important utilities. First, reading out the ancilla state enables quantum logic spectroscopy of the signal field $\Delta$. Second, conditioned on the ancilla measurement result, the clock atom is projected to a known state in the measurement basis, which enables repeated interrogation. We verify this by measuring the clock qubit state in the $x$-basis via a $\pi/2$ pulse with a variable phase $\phi$ before detection (Fig. 4c). We find the clock qubit is projected onto the |+ket|$\rangle$| + ⟩ (|ket|$\rangle$| - ⟩) state conditioned on the ancilla measurement result $|0\rangle$ ($|1\rangle$) in the $z$-basis, with a contrast of $0.84(4)$ ($0.73(4)$). Erasure excision 57, 18 is applied to discard shots where shelving has failed (Methods).
@@
Figure 5: Ancilla parity readout and measurement-based Bell state generation. a, An ancilla qubit (middle) is prepared in parallel with a Bell pair, which is then used in Ramsey interrogation. A parity signal is measured either directly on the Bell pair or through ancilla-based parity readout. b, Direct readout of the Bell pair parity signal (pink) is compared with a single ancilla readout (orange) yielding a weight-2 parity signal oscillating at the same frequency and phase within error bars (Methods). c, A circuit for measurement-based preparation of a long-range Bell pair, composed of 3-qubit cluster state generation followed by MCR on the middle ancilla qubit. Post-selection on the ancilla result $|0\rangle$ projects the two exterior atoms to the Bell state $|\Phi^{+}\rangle$. d, Characterization of the resulting Bell state with population overlap of $0.71(2)$ in the even parity subspace $\{|00\rangle,|11\rangle\}$. The parity oscillation of the resulting state yields a contrast of $0.52(3)$, resulting in an overall raw Bell state fidelity of $0.61(2)$. Erasure excision is applied to discard shots where shelving has failed. e, A proposal for a continuous clock based on a modular architecture (see Outlook for details).
@@
Ideally, non-destructive readout enables repetitive measurements of the same qubit or same register. This can be achieved by resetting the ancilla qubit and repeating the readout cycle. In our case, we utilize an 18 $\mu$s fast imaging method 57 which is much shorter than the qubit coherence time but results in a high probability of losing the ancilla atom during imaging. We thus realize repeated rounds of measurement and reset by replacing the ancilla atom with a reserve atom, which was shelved during ancilla MCR, using dynamical reconfiguration. This yields a dead time of only 2.9 ms dominated by shelving and unshelving operations. Alternatively, we demonstrate reusing the same ancilla atom by applying a 10-ms low-loss imaging followed by re-cooling (Methods, Ext. Data Fig. 10).
@@
We utilize ancilla-based readout to perform fast and repeated phase detection ($\hat{X}$ measurement) cycles with clock qubit reset (Fig. 4d,e). Each round $j$ consists of a fixed evolution time $T_{j}$ followed by ancilla-based readout. Then, we perform one final sensing round with a variable time $t$, where we verify the clock qubit coherence via direct readout. In this case, we choose the evolution time $T_{j}$ such that, in the absence of a phase error, clock qubits are in the |+ket|$\rangle$| + ⟩ state. In Fig. 4e, we plot the resulting Ramsey signal for up to three consecutive measurement rounds. We use information from ancilla detection in previous rounds to remove shots where the ancilla state was toggled from $|1\rangle$ to $|0\rangle$ indicating an error has occurred. We contrast this with reference qubits that did not interact with ancillas and find an improved contrast for the ancilla-based detection.
Thus, our demonstration further shows that in the presence of noise, the coherence of the information-carrying qubit can be partially restored through ancilla-based measurement. In the metrological context, such fast detection and reset is advantageous for low dead-time clock operation, minimizing the Dick effect 8. Our scheme could also find use in optimal multi-ensemble schemes where phase information from the repeatedly measured ensemble is used to correct for phase slips in the longer evolving ensemble 5.
@@
We now generalize our experimental demonstration of ancilla-based readout to multi-qubit measurements. In a multi-qubit measurement, an ancilla is entangled with multiple qubits and then measured. For example, a parity signal from an entangled quantum register, which has enhanced sensitivity to the signal field, can be read out via a single ancilla qubit. Additionally, the measurement projects the register qubits to a well-defined parity subspace. The latter may prove especially useful for the preparation of large-scale long-range entangled states, such as GHZ states, through ancilla measurements 21, 22, 23, 24, for which we show a proof-of-principle for optical qubits. We note that multi-qubit measurements, measuring higher weight observables, are also central to stabilizer-based quantum error correction 39.
@@
We demonstrate the lowest-order version of such schemes, a weight-2 parity measurement, where the ancilla is entangled with two different clock qubits before measurement. To do this, we first prepare an input state composed of an ancilla atom in the clock state $|1\rangle$ and two atoms in a Bell state (Fig. 5a). We then perform a $XX$ parity measurement by entangling both clock qubits with the ancilla and measuring the ancilla’s state. In Fig. 5b, we compare the direct readout of a parity signal from the Bell pair (pink) with the ancilla readout (orange). We find that indeed the single ancilla atom observable now oscillates at a frequency of twice the single atom detuning $\Delta$. This frequency, as well as the phase of the signal, is consistent with a direct readout of the Bell pair parity (Methods). We note that the contrast of these signals is largely limited by the state preparation step, which in our one-dimensional architecture includes a state exchange operation (Ext. Data Fig. 11).
@@
We then turn to demonstrate another relevant application of the conditional parity projection in a weight-2 ancilla MCR: long-range entanglement induced by measurement 21, 22, 23, 24. We first prepare a 3-qubit cluster state and then perform MCR of the central qubit in the $x$-basis (Fig. 5c). As $ZXZ$ is a stabilizer of the cluster state, this measurement generates a Bell pair with a well-defined parity for the exterior atoms, set by the initial state preparation. When post-selecting on the ancilla result $|0\rangle$, we find a Bell state with population $0.71(2)$ in the even parity subspace $\{|00\rangle,|11\rangle\}$ and a parity oscillation contrast of $0.52(3)$, yielding an overall raw fidelity of $0.61(2)$ (Fig. 5d). If instead we post-select on the ancilla result $|1\rangle$, we find the population projection is the opposite, with $0.64(2)$ in the odd parity subspace $\{|01\rangle,|10\rangle\}$. The scheme demonstrated here can be considered a heralded preparation, which could be made deterministic by a feedback operation 21, 22, 23, 24. We note that with improvements in fidelity, and with sufficiently fast feedback, this could then be readily extensible to perform measurement-based preparation of large-scale GHZ states with only a constant depth circuit 22, 21, 23. This is in contrast to linear or log-depth standard unitary circuits 31, which could be instrumental for generating large-scale optical clock GHZ states in future devices.
@@
  Our results demonstrate how universal quantum processing and ancilla-based readout can be integrated with a neutral atom optical clock. Extending such schemes to larger two-dimensional tweezer clock systems 14 would enable the creation of multiple copies of GHZ states at a given size. We note in particular that the schemes demonstrated here could be easily generalized to two dimensions by performing all operations in parallel for a set of rows to generate multiple copies of GHZ states, e.g., by combining an acousto-optical deflector (AOD) with a spatial light modulator. Deviations from the magic-wavelength condition due to the multi-frequency nature of AODs 13 could be compensated via atom movements, as proposed in Ref. 16. We see this as a viable avenue towards quantum-enhanced operation of neutral atom clocks using cascades of GHZ states 1, 4 combined with classically optimal interrogation schemes and quadrature readout 5, 48, 16, 49. Importantly, mid-circuit ancilla-based detection or mid-circuit low-loss imaging (Ext. Data Fig. 10), as shown here, could be directly used to implement such classical optimal interrogation.
@@
Upgrading the clock laser system used in this work to match the current state-of-the-art 33, 14 would enable generating significantly larger GHZ states than demonstrated here (Ext. Data Fig. 6, Methods). We then expect the maximum GHZ state size will eventually be limited by a combination of clock laser phase noise and CZ Rydberg gate fidelity. In this context, we note that the CZ gate fidelity is currently dominated by phase noise of the Rydberg laser (Ext. Data Fig. 3, Methods), which could easily be improved by implementing cavity filtering techniques 58. Further, utilizing native, beyond-two-qubit gates 27, 17 could reduce the required circuit depth. Finally, GHZ states can be prepared with constant depth circuits via measurement-based schemes 21, 22, for which we showed first steps with optical clock qubits in Fig. 5c,d.
@@
Our results further point to a modular device with distinct quantum processing and sensing modules, realizing a vision of a quantum computer connected to a high-precision sensor (Fig. 5e). The processing module could be used to prepare entangled probe states in parallel to sensing, followed by a state exchange operation between the sensor and the processor. Subsequently, readout could be performed in the processing module in parallel to sensing, before preparing new entangled probe states, yielding very low dead time operation. The processor module could also perform non-trivial multi-qubit stabilizer measurements on the sensing qubits. Sensing and multi-qubit ancilla readout could then be cycled several times before further state preparation is needed 59. Options for fully realizing this vision could be based on a multi-zone approach with a single species 24, a dual-species 52 device composed of alkaline-earth and alkali atoms (e.g., strontium and rubidium), an optical-metastable-ground-state qubit architecture 55, 60 or combinations thereof. Similar capabilities could be implemented to improve precision measurements with molecules 61. Several of our results can be viewed as proof-of-principle demonstrations towards this modular vision, including state exchange operations (Ext. Data Fig. 11) and higher-weight ancilla-based readout.
More generally, our work highlights instances where entanglement can potentially improve metrological performance in hybrid devices with sensing and quantum processing capabilities. Such devices could be part of next-generation quantum technologies from field-transportable sensors 62 to state-of-the-art optical clocks 33 in the quest to push the limits of precision measurements.
Note — During the completion of this manuscript, we became aware of a related work 63.

### v2

@@
Figure 4: Repeated ancilla-based quantum logic spectroscopy (QLS). a, We consider three ensembles of atoms: reference atoms (R; red), clock atoms (C; purple), and ancillas (A; blue). An entangling gate maps the state of clock atoms to the state of ancilla atoms before mid-circuit readout (MCR) of the ancilla state. b, Logic spectroscopy via ancilla atoms. A Ramsey interrogation is performed in parallel on reference atoms (R) and clock atoms (C). A clock-ancilla CNOT gate maps the state onto ancilla atoms (A) for readout. c, Ancilla-based $\hat{X}$-measurement. A non-destructive measurement of the state $|\psi\rangle_{C}$ is realized by shelving clock qubits ($S$ operation) followed by ancilla MCR. We show that following unshelving ($S^{\dagger}$), the state of the clock qubit is projected to a well-defined state in the measurement basis, conditioned on the ancilla state. d, $\rm{N_{cyc}}$ repeated rounds of QLS with minimal (2.9 ms) dead time, enabled by ancilla replacement via array reconfiguration. Each round $j$ consists of a fixed evolution time $T_{j}$ followed by ancilla-based $\hat{X}$ measurement before a final round where the clock qubit is directly measured in a Ramsey sequence with variable time t. e, We find improved contrast in the final measurement round when conditioned on ancilla results (see text) in all previous rounds (purple markers) as compared to reference clock atoms which are protected during MCR but do not interact with ancillas (red markers).
@@
Having used entanglement to prepare metrological probe states, we now turn to use it to read out the resulting state. Specifically, we demonstrate a method based on mapping quantum information from a set of sensor qubits to a set of ancillary qubits, which are measured while the sensor qubits are quickly available for sensing once more. Such repeated, non-destructive, ancilla-based measurements, are a central building block in quantum error detection and correction 41. In a metrological context, repeated mid-evolution measurement and reset cycles are desirable for efficient multi-ensemble schemes 5. These schemes enable extended dynamic range and effective coherence time, even in the absence of entangled probes. However, fast mid-circuit measurement and reset of atomic qubits are challenging due to the overhead in resetting both the motional state and the electronic state of a qubit after direct measurement.
@@
To efficiently realize ancilla-based readout in atomic systems, information-carrying data qubits must first be protected from global imaging light and their coherence must be preserved during mid-circuit readout (MCR) of ancilla qubits. Such protection can be realized by using dual-species arrays 56, cavity-assisted readout 57, a separate readout zone 24, or shelving to states dark to imaging light 58, 59, 60. In our system, we take the last approach and have recently realized MCR 45 by shelving optical qubits into superpositions of motional states which are long-lived and dark to imaging light.
@@
By performing high-fidelity entangling gates between information-carrying (clock) qubits and ancilla qubits before MCR (Fig. 4a), we can realize a form of quantum logic spectroscopy (QLS) with neutral atom optical clock qubits, as originally introduced for mixed-species ion traps 20. In our case, we use local control to prepare certain atoms in superposition states sensitive to the clock laser detuning $\Delta$, while ancilla atoms remain in the insensitive state $|1\rangle$. Clock qubits acquire a time-dependent phase $\Delta\cdot t$ which is mapped onto population in a Ramsey sequence. A subsequent clock-ancilla CNOT gate (controlled on qubit state $|0\rangle$) then maps this signal onto the ancilla state (Fig. 4b). This sequence realizes an ancilla-based detection of $\hat{X}$ of the clock qubit.
@@
To realize a non-destructive ancilla-based $\hat{X}$-measurement, we then shelve clock qubits and read out the ancilla state. To this end, we ramp up the depth of ancilla traps by a factor of 5 (leaving traps with clock atoms unchanged) such that they are non-resonant with the global motional shelving drive. Following shelving of clock qubits, we perform ancilla MCR of the signal field $\Delta$. Conditioned on the ancilla measurement result, the clock atom is projected to a known state in the measurement basis, which enables repeated interrogation. We verify this by measuring the clock qubit state in the $\hat{X}$-basis via a $\pi/2$ pulse with a variable phase $\phi$ before detection (Fig. 4c). We find the clock qubit is projected onto the |+ket|$\rangle$| + ⟩ (|ket|$\rangle$| - ⟩) state conditioned on the ancilla measurement result $|0\rangle$ ($|1\rangle$) in the $\hat{Z}$-basis, with a contrast of $0.84(4)$ ($0.73(4)$). These results are conditioned on successful one-way shelving 45, 18 (Methods).
@@
Figure 5: Ancilla parity readout and measurement-based Bell state generation. a, An ancilla qubit (middle) is prepared in parallel with a Bell pair, which is then used in Ramsey interrogation. A parity signal is measured either directly on the Bell pair or through ancilla-based parity readout. b, Direct readout of the Bell pair parity signal (pink) is compared with a single ancilla readout (orange) yielding a weight-2 parity signal oscillating at the same frequency and phase within error bars (Methods). c, A circuit for measurement-based preparation of a long-range Bell pair, composed of 3-qubit cluster state generation followed by MCR on the middle ancilla qubit. Conditioning on the ancilla result $|0\rangle$ ($\approx 45\%$ of experimental shots) projects the two exterior atoms to the Bell state $|\Phi^{+}\rangle$. d, Characterization of the resulting Bell state with population overlap of $0.71(2)$ in the even parity subspace $\{|00\rangle,|11\rangle\}$. The parity oscillation of the resulting state yields a contrast of $0.52(3)$, resulting in an overall raw Bell state fidelity of $0.61(2)$. Data is conditioned on the successful shelving of the two exterior qubits ($\approx 82\%$ of experimental shots). e, A proposal for a continuous clock based on a modular architecture (see Outlook for details).
@@
Ideally, non-destructive readout enables repetitive measurements of the same qubit or the same register. This can be achieved by resetting the ancilla qubit and repeating the readout cycle. In our case, we utilize an 18 $\mu$s fast imaging method 61 which is much shorter than the qubit coherence time but results in a high probability of losing the ancilla atom during imaging. We thus realize repeated rounds of measurement and reset by replacing the ancilla atom with a reserve atom, which was shelved during ancilla MCR, using dynamical reconfiguration. This yields a dead time of only 2.9 ms dominated by shelving and unshelving operations. Alternatively, we demonstrate reusing the same ancilla atom by applying a 10-ms low-loss imaging followed by re-cooling (Methods, Ext. Data Fig. 10).
@@
We utilize ancilla-based readout to perform fast and repeated phase detection ($\hat{X}$-measurement) cycles with clock qubit reset (Fig. 4d). In Fig. 4e, we plot the resulting Ramsey signal for up to three consecutive measurement rounds. We use information from ancilla detection in previous rounds to remove shots where the ancilla state was toggled from $|1\rangle$ to $|0\rangle$ (maintaining $78\%$ of the experimental shots for a single round of ancilla detection), indicating a phase flip. We contrast this with reference qubits that did not interact with ancillas and find an improved contrast for the ancilla-based detection. For the latter, the contrast is now limited by the fidelity of shelving and unshelving operations as part of the MCR.
While here ancillas are used to measure a single qubit, extending our demonstration to a larger ensemble of ancillas interacting with an ensemble of clock qubits would enable single-shot phase estimation of the global field $\Delta$. Information on phase deviation outside of the dynamic range could be used to correct the phase estimation of the longer evolving ensemble realized by the reference qubits in our demonstration 5. Such fast detection and reset are further advantageous for low dead-time clock operation, reducing the instability due to the Dick effect 8. On the other hand, using ancillas in a single-species array comes at the expense of a reduced number of clock qubits, increasing quantum projection noise, such that the metrological gain depends on the balance of the two.
@@
We now generalize our experimental demonstration of ancilla-based readout to multi-qubit measurements. In a multi-qubit measurement, an ancilla is entangled with multiple qubits and then measured. The measurement projects the register qubits to a well-defined parity subspace. This may prove especially useful for the preparation of large-scale long-range entangled states, such as GHZ states, through ancilla measurements 21, 22, 23, 24, for which we show a proof-of-principle for optical qubits. We note that multi-qubit measurements, measuring higher weight observables, are also central to stabilizer-based quantum error correction 41.
@@
We demonstrate the lowest-order version of such schemes, a weight-2 parity measurement, where the ancilla is entangled with two different clock qubits before measurement. To do this, we prepare an input state composed of an ancilla atom in the clock state $|1\rangle$ and two atoms in a Bell state and perform a $\hat{X}\hat{X}$ parity measurement (Fig. 5a). In Fig. 5b, we compare the direct readout of a parity signal from the Bell pair (pink) with the ancilla readout (orange). We find that indeed the single ancilla atom observable now oscillates at a frequency of twice the single atom detuning $\Delta$. This frequency, as well as the phase of the signal, is consistent with a direct readout of the Bell pair parity (Methods).
@@
We now turn to demonstrate another relevant application of the conditional parity projection in a weight-2 ancilla MCR: long-range entanglement induced by measurement 21, 22, 23, 24. We first prepare a 3-qubit cluster state and then perform MCR of the central qubit in the $\hat{X}$-basis (Fig. 5c). As $\hat{Z}\hat{X}\hat{Z}$ is a stabilizer of the cluster state, this measurement generates a Bell pair with a well-defined parity for the exterior atoms. Conditioning on the ancilla result $|0\rangle$, we find a Bell state with population $0.71(2)$ in the even parity subspace $\{|00\rangle,|11\rangle\}$ and a parity oscillation contrast of $0.52(3)$, yielding an overall raw fidelity of $0.61(2)$ (Fig. 5d). If instead we condition on the ancilla result $|1\rangle$, we find the state is projected to the odd parity Bell state, which is not useful in a metrological context, with $0.64(2)$ in the odd parity subspace $\{|01\rangle,|10\rangle\}$. The scheme demonstrated here can be considered a heralded preparation, which could be made deterministic by a feedback operation 21, 22, 23, 24. We note that with improvements in fidelity, and with sufficiently fast feedback, this could then be readily extended to perform measurement-based preparation of large-scale GHZ states with only a constant depth circuit 22, 21, 23. This is in contrast to linear or log-depth standard unitary circuits 31, which could be instrumental for generating large-scale optical clock GHZ states in future devices.
@@
  Our results demonstrate how universal quantum processing and ancilla-based readout can be integrated with a neutral atom optical clock. Extending such schemes to larger two-dimensional tweezer clock systems 14 would enable the creation of multiple copies of GHZ states at a given size. We note in particular that the schemes demonstrated here could be easily generalized to two dimensions by performing all operations in parallel for a set of rows to generate multiple copies of GHZ states. We see this as a viable avenue towards quantum-enhanced operation of neutral atom clocks using cascades of GHZ states 1, 4 combined with classically optimal interrogation schemes and quadrature readout 5, 52, 16, 53.
@@
We have demonstrated a CZ fidelity of $0.9962(4)$ , averaged over the two-qubit symmetric stabilizer states, setting a new state-of-the-art for entangling gate fidelities in long-lived states of neutral atoms, as well as the full toolbox needed to execute relevant algorithms for metrology and computation with optical clock qubits. The main limitation to current circuit fidelities is the clock laser phase noise resulting in dephasing during circuit idle times. The ancilla-based MCR is further limited by the fidelity of shelving qubits in motional states. Upgrading the clock laser system used in this work to match the current state-of-the-art 35, 14 would enable generating significantly larger GHZ states than demonstrated here (Ext. Data Fig. 6, Methods). We then expect the maximum GHZ state size will eventually be limited by a combination of clock laser phase noise and CZ gate fidelity. In this context, we note that the CZ gate fidelity is currently dominated by phase noise of the Rydberg laser (Ext. Data Fig. 4, Methods), which could easily be improved by implementing cavity filtering techniques 62. Further, utilizing native, multi-qubit gates 63, 17 or effective all-to-all connectivity 19 could reduce the required circuit depth. Ancilla-based MCR would also improve with a higher stability clock laser and better control and homogenization of trap waists 45. Finally, GHZ states can be prepared with constant depth circuits via measurement-based schemes 21, 22, for which we showed first steps with optical clock qubits in Fig. 5c,d.
@@
Our results further point to a modular device with distinct quantum processing and sensing modules, realizing a vision of a quantum computer connected to a high-precision sensor (Fig. 5e). The processing module could be used to prepare entangled probe states in parallel to sensing, followed by a state exchange operation between the sensor and the processor. Subsequently, readout could be performed in the processing module in parallel to sensing, before preparing new entangled probe states, yielding very low dead time operation. Options for fully realizing this vision could be based on a multi-zone approach with a single species 24, a dual-species 56 device composed of alkaline-earth and alkali atoms (e.g., strontium and rubidium), an optical-metastable-ground-state qubit architecture 59, 64 or combinations thereof. Several of our results can be viewed as proof-of-principle demonstrations towards this modular vision, including state exchange operations (Ext. Data Fig. 11) and higher-weight ancilla-based readout. Such a device could be part of next-generation quantum technologies from field-transportable sensors to state-of-the-art optical clocks 35 in the quest to push the limits of precision measurements.